Let a and b be any real numbers. Define a*b to be 2a+b. Which of the following is the largest?

1*(2*(3*4))
((1*2)*3)*4
(1*(2*3))*4
They are equal

Answer by stanbon(75887)

Let a and b be any real numbers. Define a*b to be 2a+b. Which of the following is the largest?
-------------
1*(2*(3*4))
= 1*(6+4) = 1*10 = 2+10 = 12
-----------------------------
((1*2)*3)*4
= ((4)*3)*4
= (13*4)
= 30
===========
(1*(2*3))*4
= (1*(7)*4)
= 9*4
= 22
================
Cheers,
Stan H.
=============
They are equal